# MoPD: Mixture-of-Prompts Distillation for Vision-Language Models

### v1

@@ Abstract
Soft prompt learning methods are effective for adapting vision-language models (VLMs) to downstream tasks. Nevertheless, empirical evidence reveals a tendency of existing methods that they overfit seen classes and exhibit degraded performance on unseen classes. This limitation is due to the inherent bias in the training data towards the seen classes. To address this issue, we propose a novel soft prompt learning method, named Mixture-of-Prompts Distillation (MoPD), which can effectively transfer useful knowledge from hard prompts manually hand-crafted (a.k.a. teacher prompts) to the learnable soft prompt (a.k.a. student prompt), thereby enhancing the generalization ability of soft prompts on unseen classes. Moreover, the proposed MoPD method utilizes a gating network that learns to select hard prompts used for prompt distillation. Extensive experiments demonstrate that the proposed MoPD method outperforms state-of-the-art baselines especially on on unseen classes.
@@ I Introduction
Recently, some works (e.g., ProGrad [13] and KgCoOp [14]) have shown that soft prompt learning methods tend to overfit the base (seen) classes and generalize poorly on the new (unseen) classes, while VLMs with hard prompts (e.g., a photo of a [CLASS], where [CLASS] denotes the placeholder for the class name) can achieve good zero-shot recognition performance on new classes. To utilize the useful knowledge captured in the pre-trained VLMs through hard prompts, ProGrad regulates the gradient direction to mitigate potential conflicts between hard prompts and soft prompts. Similarly, KgCoOp mitigates the risk of knowledge loss by minimizing the embedding discrepancy between soft prompts and hard prompts.
However, the aforementioned methods solely rely on a single hard prompt, without considering the use of multiple hard prompts to enhance the learning of soft prompts. In fact, compared with a single hard prompt, a pool with multiple hard prompts may provide more useful knowledge for prompt-based models and such a pool is commonly available. Besides, knowledge distillation techniques [15, 16, 17, 18] are an effective approach to transfer knowledge, but they are unexplored in existing soft prompt learning methods for VLMs.
@@
To solve these issues, we propose a Mixture-of-Prompts Distillation (MoPD) method. Specifically, the proposed MoPD method utilizes a trainable gating network to select one or more instance-specific hard prompts from a pool of hard prompts, guiding the learning of the soft prompt, where the gating network assigns weights to teacher prompts based on image features during the prompt learning stage. Extensive experiments on 11 datasets demonstrate that the proposed MoPD method outperforms state-of-the-art baseline methods under several settings (i.e., base-to-new, few-shot classification, and domain generalization).
@@ II-C Soft Prompt Learning
To solve the above issue in prompting, soft prompt learning methods are proposed to learn a group of prompting vectors while fixing parameters in the backbone. For example, CoOp [11] utilizes learnable prompting vectors to replace the prompting words like “a photo of a” and adapts to downstream tasks. However, CoOp tends to overfit seen classes, thereby limiting its ability to generalize to unseen classes within the same dataset. To address the issue of weak generalization ability, CoCoOp [12] trains a lightweight neural network to generate input-conditional vectors for each image, collaborating with prompting vectors to dynamically generate prompting vectors. As discussed before, ProGrad and KgCoOp aim to use a hard prompt to guide the learning of the soft prompt in different ways.
@@ III Methodology
Figure 2: Overview of the proposed MoPD method.
@@ III-A Preliminaries
Given a set of class names with a total number of $C$ classes, a hard prompt like “a photo of a [$\mathrm{CLASS}$]” is inputted into the text encoder to generate textual embedding $\boldsymbol{t}_{\mathrm{hard}}^{c}$ for the $c$-th class. The visual embedding $\boldsymbol{f}$ is extracted by the image encoder for an image $\boldsymbol{x}$. The prediction probability of image $\boldsymbol{x}$ belonging to class $y$ is formulated as
@@
where $\tau$ is a temperature parameter and $\cos\left(\cdot,\cdot\right)$ denotes the cosine similarity.
However, hard prompts (e.g., “a photo of a [$\mathrm{CLASS}$]”), which are straightforward for humans, may not be optimal for VLMs. To enhance the adaptability of prompts for various downstream tasks, the CoOp method replaces such prompting words with learnable prompting vectors, which are called soft prompts. This soft prompting approach has demonstrated improved performance on multiple downstream datasets compared with using hard prompts [1, 19]. Specifically, the soft prompt consists of learnable prompting vectors of length $M$, where each prompting vector has the same dimension as the word embedding in VLMs, and a class name. Hence, the soft prompt can be defined as $\left\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{M},\left[\mathrm{CLASS}\right]\right\}$. The textual embedding of $c$-th class can be obtained by feeding the soft prompt into the text encoder, and it is denoted by $\boldsymbol{t}_{\mathrm{soft}}^{c}$. Then the prediction probability of image $\boldsymbol{x}$ belonging to class $y$ is formulated as
$p_{\mathrm{soft}}\left(y|\boldsymbol{x}\right)=\frac{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{soft}}^{y},\boldsymbol{f}\right)/\tau\right)}{\sum\nolimits_{c=1}^{C}{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{soft}}^{c},\boldsymbol{f}\right)/\tau\right)}}.$ (2)
The soft prompt can be learned by minimizing the cross-entropy loss as
@@ III-B2 Mixture-of-Prompts Distillation (MoPD)
Usually there may exist significant variations among images within the same dataset [24] and hence a single hard prompt is insufficient to accommodate the preferences of all images. To alleviate this issue, it is intuitive to utilize multiple hard prompts to fit such preferences and a straightforward way is to randomly select hard prompts from a hard prompt pool. However, this method still does not take the individual preferences of each image into consideration and cannot filter out potentially noisy hard prompts from the pool.
@@ IV Experiments
In this section, we evaluate MoPD on three settings: (1) base-to-new generalization, (2) few-shot classification, (3) domain generalization. We also analyze the effect of each module and parameter through ablation study and parameter analysis. We further analyze the impact of noisy prompts.
@@
Following [11, 12, 13, 14], the base-to-new generalization and few-shot image classification are conducted on 11 image classification datasets, including ImageNet [26] and Caltech101 [27] for generic object classification, OxfordPets [28], StanfordCars [29], Flowers102 [30], Food101 [31], and FGVCAircraft [32] for fine-grained visual categorization, SUN397 [33] for scene recognition, DTD [34] for texture classification, EuroSAT [35] for satellite image classification, and UCF101 [36] for action recognition. Besides, the domain generalization is conducted on ImageNet and its variants, including ImageNetV2 [37], ImageNet-Sketch [38], ImageNet-A [39], and ImageNet-R [40].
@@
We compare with several state-of-the-art soft prompt learning methods, including CoOp [11], CoCoOp [12], ProGrad [13], and KgCoOp [14]. CLIP [1] is also in comparison to compare with the generalization performance of hard prompts.
@@ IV-A Results on Base-to-New Generalization
We follow the base-to-new generalization setting of CoCoOp, where all classes are split into disjointed base and new groups and all the methods in comparison are trained on the base classes and evaluated on base and new classes. The experimental results shown in Table I reveal the following observations.
Dataset Set CLIP CoOp CoCoOp ProGrad KgCoOp MoPD
Average Base 69.34 82.64 80.47 82.48 80.73 81.40
New 74.22 68.00 71.69 70.72 73.61 74.69
H 71.70 74.61 75.83 76.15 77.00 77.90
ImageNet Base 72.43 76.46 75.98 77.02 75.83 76.87
New 68.14 66.31 70.43 66.66 69.96 70.43
H 70.22 71.02 73.10 71.46 72.78 73.51
Caltech 101 Base 96.84 98.11 97.96 98.02 97.72 98.07
New 94.00 93.52 93.81 93.89 94.39 94.90
H 95.40 95.76 95.84 95.91 96.03 96.46
Oxford Pets Base 91.17 94.24 95.20 95.07 94.65 95.17
New 97.26 96.66 97.67 97.63 97.76 97.47
H 94.12 95.43 96.43 96.33 96.18 96.31
Stanford Cars Base 63.37 76.20 70.49 77.68 71.76 75.63
New 74.89 69.14 73.59 68.63 75.04 73.10
H 68.65 72.49 72.01 72.88 73.36 74.34
Flowers 102 Base 72.08 97.63 94.87 95.54 95.00 95.80
New 77.80 69.55 71.75 71.87 74.73 74.43
H 74.83 81.23 81.71 82.03 83.65 83.77
Food101 Base 90.10 89.44 90.70 90.37 90.50 90.77
New 91.22 87.50 91.29 89.59 91.70 91.67
H 90.66 88.46 90.99 89.98 91.09 91.22
FGVC Aircraft Base 27.19 39.24 33.41 40.54 36.21 37.43
New 36.29 30.49 23.71 27.57 33.55 35.13
H 31.09 34.30 27.74 32.82 34.83 36.24
SUN397 Base 69.36 80.85 79.74 81.26 80.29 81.13
New 75.35 68.34 76.86 74.17 76.53 76.47
H 72.23 74.07 78.27 77.55 78.36 78.73
DTD Base 53.24 80.17 77.01 77.35 77.55 77.27
New 59.90 47.54 56.00 52.35 54.99 57.47
H 56.37 59.68 64.85 62.45 64.35 65.92
EuroSAT Base 56.48 91.54 87.49 90.11 85.64 85.77
New 64.05 54.44 60.04 60.89 64.34 72.50
H 60.03 68.27 71.21 72.67 73.48 78.58
UCF101 Base 70.53 85.14 82.33 84.33 82.89 81.53
New 77.50 64.47 73.45 74.94 76.67 78.07
H 73.85 73.37 77.64 79.35 79.65 79.76
@@
    In terms of the harmonic mean accuracy, MoPD outperforms all baseline methods, showcasing its excellent generalization capability on the base-to-new setting. Specifically, MoPD attains the highest harmonic mean accuracy in 10 out of 11 datasets.
@@
    In terms of the average performance on the new classes, MoPD surpasses all baseline methods. MoPD exhibits a significant performance improvement over soft prompt learning methods and outperforms CLIP on 7 out of 11 datasets. This demonstrates that the selected teacher prompts can effectively guide the soft prompt in acquiring useful knowledge and enhancing its generalization capability.
@@
    In terms of the average performance on the base classes, CoOp and ProGrad perform better among all methods since they may overfit the base classes, leading to poor performance on new classes. MoPD performs slightly inferior to CoOp and ProGrad on the base classes, but substantially outperforms them on the new classes, while MoPD outperforms CoCoOp and KgCoOp on both the base and new classes.

### v2

@@ Abstract
Soft prompt learning methods are effective for adapting vision-language models (VLMs) to downstream tasks. Nevertheless, empirical evidence reveals that existing methods tend to overfit seen classes and exhibit degraded performance on unseen classes. This limitation is due to the inherent bias in the training data towards the seen classes. To address this issue, we propose a novel soft prompt learning method, named Mixture-of-Prompts Distillation (MoPD), which can effectively transfer useful knowledge from hard prompts manually hand-crafted (a.k.a. teacher prompts) to the learnable soft prompt (a.k.a. student prompt), thereby enhancing the generalization ability of soft prompts on unseen classes. Moreover, the proposed MoPD method utilizes a gating network that learns to select hard prompts used for prompt distillation. Extensive experiments demonstrate that the proposed MoPD method outperforms state-of-the-art baselines, especially on unseen classes.
@@ I Introduction
Recently, some works (e.g., ProGrad [13] and KgCoOp [14]) have shown that soft prompt learning methods tend to overfit the base (seen) classes and generalize poorly on the new (unseen) classes, while VLMs with hard prompts (e.g., a photo of a [CLASS], where [CLASS] denotes the placeholder for the class name) can achieve good zero-shot recognition performance on new classes. To utilize the useful knowledge captured in the pre-trained VLMs through hard prompts, ProGrad regulates the gradient direction to mitigate potential conflicts between hard prompts and soft prompts. KgCoOp mitigates the risk of knowledge loss by minimizing the embedding discrepancy between soft prompts and hard prompts. Similarly, to align with pre-trained CLIP’s generalization, LASP [15] maximizes the probability of the learned prompts being correctly classified with respect to pre-defined hard prompts.
However, the aforementioned methods solely rely on a single hard prompt, without considering the use of multiple hard prompts to enhance the learning of soft prompts. In fact, compared with a single hard prompt, a pool with multiple hard prompts may provide more useful knowledge for prompt-based models, and such a pool is commonly available. Besides, knowledge distillation techniques [16, 17, 18, 19] are an effective approach to transfer knowledge, but they are unexplored in existing soft prompt learning methods for VLMs.
@@
To solve these issues, we propose a Mixture-of-Prompts Distillation (MoPD) method. Specifically, the proposed MoPD method utilizes a trainable gating network to select one or more instance-specific hard prompts from a pool of hard prompts, guiding the learning of the soft prompt, where the gating network assigns weights to teacher prompts based on image features during the prompt learning stage. Extensive experiments on 11 datasets demonstrate that the proposed MoPD method outperforms state-of-the-art baseline methods under several settings (i.e., base-to-new, few-shot classification, domain generalization, and cross-dataset evaluation).
@@ II-C Soft Prompt Learning
To solve the above issue in prompting, soft prompt learning methods are proposed to learn a group of prompting vectors while fixing parameters in the backbone. For example, CoOp [11] utilizes learnable prompting vectors to replace the prompting words like “a photo of a” and adapts to downstream tasks. SgVA-CLIP [25] combines prompt learning and adapter [26], producing more discriminative visual information by comprehensively considering both adapted visual feature space and pre-trained cross-modal feature space, thus enhancing few-shot image classification. However, the previous methods tend to overfit seen classes, thereby limiting their ability to generalize to unseen classes within the same dataset. To address the issue of weak generalization ability, CoCoOp [12] trains a lightweight neural network to generate input-conditional vectors for each image, collaborating with prompting vectors to dynamically generate prompting vectors. As discussed before, ProGrad, KgCoOp, and LASP aim to use a hard prompt to guide the learning of the soft prompt in different ways. However, using only textual prompt learning as the aforementioned methods did is sub-optimal. To alleviate that, MaPLe [27] introduces a branch-aware multi-modal prompt learning method that simultaneously tunes vision and language branches together by sharing soft prompts across both modalities. Unlike MaPLe, PromptSRC [28] trains independent soft prompts for text and image encoders, and introduces a self-regularization framework to learn general features from the frozen CLIP. Similarly, CoPrompt [29] employs consistency constraints for prompt learning to enhance the generalization and adopts a large language model to generate more descriptive sentences. MetaPrompt [30] enhances the generalization by utilizing gradients from meta-test subtasks to regularize soft prompts.
@@ III Methodology
Figure 2: Overview of the proposed MoPD method, where solid arrows represent forward propagation processes and dashed arrows indicate backward propagation processes.
@@ III-A Preliminaries
Given a set of class names with a total number of $C$ classes, a hard prompt like “a photo of a $\mathrm{CLASS}$]” is first tokenized into discrete tokens and converted into word embeddings. These embeddings are then fed into the text encoder to generate textual embedding $\boldsymbol{t}_{\mathrm{hard}}\in\mathbb{R}^{C\times d}$, where $d$ is the dimension (i.e., 512 for CLIP). The visual embedding $\boldsymbol{f}$ is extracted by the image encoder for an image $\boldsymbol{x}$. The prediction probability of image $\boldsymbol{x}$ belonging to the $y$-th class is formulated as
@@
where $\tau$ is a temperature parameter, $\cos\left(\cdot,\cdot\right)$ denotes the cosine similarity, and $\boldsymbol{t}_{\mathrm{hard}}^{y}$ denotes the textual embedding of the $y$-th class generated by the hard prompt.
However, hard prompts (e.g., “a photo of a $\mathrm{CLASS}$]”), which are straightforward for humans, may not be optimal for VLMs. To enhance the adaptability of prompts for various downstream tasks, the CoOp method replaces such prompting words with learnable prompting vectors, which are called soft prompts. This soft prompting approach has demonstrated improved performance on multiple downstream datasets compared with using hard prompts [1, 20]. Specifically, the soft prompt consists of learnable prompting vectors of length $M$, where each prompting vector has the same dimension as the word embedding in VLMs, and a class name. Hence, the soft prompt can be defined as $\left\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{M},\left[\mathrm{CLASS}\right]\right\}$. The textual embedding can be obtained by feeding the soft prompt into the text encoder, and it is denoted by $\boldsymbol{t}_{\mathrm{soft}}\in\mathbb{R}^{C\times d}$. Then the prediction probability of image $\boldsymbol{x}$ belonging to the $y$-th class is formulated as
$p_{\mathrm{soft}}\left(y|\boldsymbol{x}\right)=\frac{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{soft}}^{y},\boldsymbol{f}\right)/\tau\right)}{\sum\nolimits_{c=1}^{C}{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{soft}}^{c},\boldsymbol{f}\right)/\tau\right)}}{\color[rgb]{0,0,0},}$ (2)
where $\boldsymbol{t}_{\mathrm{soft}}^{y}$ denotes the textual embedding of the $y$-th class generated by the learnable soft prompt. The soft prompt can be learned by minimizing the cross-entropy loss as
@@ III-B2 Mixture-of-Prompts Distillation (MoPD)
Usually there may exist significant variations among images within the same dataset [31], and hence a single hard prompt is insufficient to accommodate the preferences of all images. To alleviate this issue, it is intuitive to utilize multiple hard prompts to fit such preferences, and a straightforward way is to randomly select hard prompts from a hard prompt pool. However, this method still does not take the individual preferences of each image into consideration and cannot filter out potentially noisy hard prompts from the pool.
@@
III-C Discussion
To see why using multiple hard prompts could be effective, we can give an analysis from the perspective of mixture-of-experts (MoE).
From a technical perspective, the gating mechanism in MoPD is analogous to the MoE paradigm [32]. In MoE, a gating network dynamically routes inputs to specialized experts based on input features, while in MoPD, the gating network acts as a “prompt router” that selects instance-specific hard prompts, which act similarly to experts in MoE, based on image features.
From the theoretical perspective, since the gating mechanism of MoPD is grounded in MoE, the theoretical principles of MoE could be applicable to MoPD. Theorem 4.1 in [33] states that “single expert performs poorly”, while Theorem 4.2 in [33] demonstrates that a router can learn cluster-center features, helping divide a complex problem into simpler sub-problems that individual experts can conquer. [33] combines Theorem 4.1 and Theorem 4.2 to conclude that “an MoE provably outperforms a single expert”. Given the connection between MoPD and MoE, those theorems provide additional theoretical support for the proposed MoPD method that could outperform SiPD.
@@ IV Experiments
In this section, we evaluate MoPD on four settings: (1) base-to-new generalization, (2) few-shot classification, (3) domain generalization, (4) cross-dataset evaluation. We also analyze the effect of each module and parameter through ablation study and parameter analysis. We further analyze the impact of noisy prompts.
@@
Following [11, 12, 13, 14], the base-to-new generalization, cross-dataset evaluation, and few-shot image classification are conducted on 11 image classification datasets, including ImageNet [34] and Caltech101 [35] for generic object classification, OxfordPets [36], StanfordCars [37], Flowers102 [38], Food101 [39], and FGVCAircraft [40] for fine-grained visual categorization, SUN397 [41] for scene recognition, DTD [42] for texture classification, EuroSAT [43] for satellite image classification, and UCF101 [44] for action recognition. Besides, the domain generalization is conducted on ImageNet and its variants, including ImageNetV2 [45], ImageNet-Sketch [46], ImageNet-A [47], and ImageNet-R [48].
@@
We compare with several state-of-the-art soft prompt learning methods, including both textual prompt (TP) learning methods (e.g., CoOp [11], CoCoOp [12], ProGrad [13], and KgCoOp [14]) and multi-modal prompt (MMP) learning methods (e.g., MaPLe [27] and MetaPrompt [30]). Hand-crafted prompt (HP) method (i.e., CLIP [1]) is also included in comparison to evaluate the generalization performance of hard prompts.
Since MoPD is optimized using traditional textual prompt learning, it has fewer learnable parameters than multi-modal prompt learning methods (e.g., MaPLe contains 600+ times more learnable parameters than MoPD). Given those fundamental architectural differences, direct comparisons would be methodologically unsuitable. For a fair comparison with MMP methods, we extend the MoPD method into an independent vision-language prompting (IVLP) version named MoPD-MMP, which learns hierarchical prompts for both the text and image encoders separately as [28], [50] did.
Methods Base New H
HP CLIP 69.34 74.22 71.70
TP CoOp 82.64 68.00 74.61
CoCoOp 80.47 71.69 75.83
ProGrad 82.48 70.72 76.15
KgCoOp 80.73 73.61 77.01
MoPD 81.40 74.69 77.90
MMP MaPLe 81.95 75.14 78.40
MetaPrompt 82.73 75.21 78.79
MoPD-MMP 83.43 75.41 79.22
(a) Average over 11 datasets
Methods Base New H
HP CLIP 72.43 68.14 70.22
TP CoOp 76.46 66.31 71.02
CoCoOp 75.98 70.43 73.10
ProGrad 77.02 66.66 71.47
KgCoOp 75.83 69.96 72.78
MoPD 76.87 70.43 73.51
MMP MaPLe 76.66 70.54 73.47
MetaPrompt 77.10 70.90 73.87
MoPD-MMP 77.60 69.91 73.55
(b) ImageNet
Methods Base New H
HP CLIP 96.84 94.00 95.40
TP CoOp 98.11 93.52 95.76
CoCoOp 97.96 93.81 95.84
ProGrad 98.02 93.89 95.91
KgCoOp 97.72 94.39 96.03
MoPD 98.07 94.90 96.46
MMP MaPLe 97.74 94.36 96.02
MetaPrompt 98.27 94.10 96.14
MoPD-MMP 98.36 94.18 96.22
(c) Caltech101
Methods Base New H
HP CLIP 91.17 97.26 94.12
TP CoOp 94.24 96.66 95.43
CoCoOp 95.20 97.67 96.42
ProGrad 95.07 97.63 96.33
KgCoOp 94.65 97.76 96.18
MoPD 95.17 97.47 96.31
MMP MaPLe 95.43 97.76 96.58
MetaPrompt 95.57 96.87 96.22
MoPD-MMP 95.43 97.78 96.59
(d) OxfordPets
Methods Base New H
HP CLIP 63.37 74.89 68.65
TP CoOp 76.20 69.14 72.50
CoCoOp 70.49 73.59 72.01
ProGrad 77.68 68.63 72.88
KgCoOp 71.76 75.04 73.36
MoPD 75.63 73.10 74.34
MMP MaPLe 72.94 74.00 73.47
MetaPrompt 75.47 74.67 75.07
MoPD-MMP 79.75 73.06 76.26
(e) StanfordCars
Methods Base New H
HP CLIP 72.08 77.80 74.83
TP CoOp 97.63 69.55 81.23
CoCoOp 94.87 71.75 81.71
ProGrad 95.54 71.87 82.03
KgCoOp 95.00 74.73 83.65
MoPD 95.80 74.43 83.77
MMP MaPLe 95.92 72.46 82.56
MetaPrompt 97.20 74.13 84.11
MoPD-MMP 97.28 75.18 84.81
(f) Flowers102
Methods Base New H
HP CLIP 90.10 91.22 90.66
TP CoOp 89.44 87.50 88.46
CoCoOp 90.70 91.29 90.99
ProGrad 90.37 89.59 89.98
KgCoOp 90.50 91.70 91.10
MoPD 90.77 91.67 91.22
MMP MaPLe 90.71 92.05 91.38
MetaPrompt 90.70 91.67 91.18
MoPD-MMP 90.85 91.75 91.30
(g) Food101
Methods Base New H
HP CLIP 27.19 36.29 31.09
TP CoOp 39.24 30.49 34.32
CoCoOp 33.41 23.71 27.74
ProGrad 40.54 27.57 32.82
KgCoOp 36.21 33.55 34.83
MoPD 37.43 35.13 36.24
MMP MaPLe 37.44 35.61 36.50
MetaPrompt 38.50 36.60 37.53
MoPD-MMP 41.18 35.55 38.16
(h) FGVCAircraft
Methods Base New H
HP CLIP 69.36 75.35 72.23
TP CoOp 80.85 68.34 74.07
CoCoOp 79.74 76.86 78.27
ProGrad 81.26 74.17 77.55
KgCoOp 80.29 76.53 78.36
MoPD 81.13 76.47 78.73
MMP MaPLe 80.82 78.70 79.75
MetaPrompt 81.80 78.50 80.12
MoPD-MMP 82.20 78.54 80.33
(i) SUN397
Methods Base New H
HP CLIP 53.24 59.90 56.37
TP CoOp 80.17 47.54 59.69
CoCoOp 77.01 56.00 64.85
ProGrad 77.35 52.35 62.44
KgCoOp 77.55 54.99 64.35
MoPD 77.27 57.47 65.92
MMP MaPLe 80.36 59.18 68.16
MetaPrompt 82.20 58.97 68.67
MoPD-MMP 81.40 57.65 67.50
(j) DTD
Methods Base New H
HP CLIP 56.48 64.05 60.03
TP CoOp 91.54 54.44 68.28
CoCoOp 87.49 60.04 71.21
ProGrad 90.11 60.89 72.67
KgCoOp 85.64 64.34 73.48
MoPD 85.77 72.50 78.58
MMP MaPLe 90.47 73.23 80.94
MetaPrompt 88.77 72.50 79.81
MoPD-MMP 86.79 78.20 82.27
(k) EuroSAT
Methods Base New H
HP CLIP 70.53 77.50 73.85
TP CoOp 85.14 64.47 73.38
CoCoOp 82.33 73.45 77.64
ProGrad 84.33 74.94 79.36
KgCoOp 82.89 76.67 79.66
MoPD 81.53 78.07 79.76
MMP MaPLe 83.00 78.66 80.77
MetaPrompt 84.47 78.40 81.32
MoPD-MMP 86.85 77.68 82.01
(l) UCF101
@@ IV-A Results on Base-to-New Generalization
We follow the base-to-new generalization setting of CoCoOp, where all classes are split into disjointed base and new groups and all the methods in comparison are trained on the base classes and evaluated on base and new classes. The experimental results shown in Table Il reveal the following observations.
@@
    In terms of the average harmonic mean accuracy, MoPD outperforms all TP and HP baseline methods, showcasing its excellent generalization capability on the base-to-new setting. Specifically, MoPD attains the highest harmonic mean accuracy in 10 out of 11 datasets.
@@
    In terms of the average performance on the new classes, MoPD surpasses all TP and HP baseline methods. MoPD exhibits a significant performance improvement over soft prompt learning methods and outperforms CLIP on 7 out of 11 datasets. This demonstrates that the selected teacher prompts can effectively guide the soft prompt in acquiring useful knowledge and enhancing its generalization capability.
@@
    In terms of the average performance on the base classes, CoOp and ProGrad perform better among all methods (except MetaPrompt and MoPD-MMP) since they may overfit the base classes, leading to poor performance on new classes. MoPD performs slightly inferior to CoOp and ProGrad on the base classes, but substantially outperforms them on the new classes, while MoPD outperforms CoCoOp and KgCoOp on both the base and new classes.
  4. 4.
    Compared with the HP and TP baseline methods, the MMP methods can achieve a better trade-off on the performance between base and new classes. In other words, the MMP methods outperform the HP and TP methods in terms of the average harmonic mean accuracy. Compared with all baseline methods including the MMP methods (i.e., MaPLe and MetaPrompt), MoPD-MMP achieves the best average performance on base classes, new classes, and harmonic mean accuracy, which demonstrates the effectiveness of the proposed MoPD method and its good compatibility with multi-modal prompt learning.